OpenStudy (anonymous):

Can anyone tell me where to Find Science? Just normal Science I Need to know If 400 Pa of pressure is applied to an area of 55 m2, which is the resulting force? Select one of the options below as your answer: A. 345 N B. 7.27 N C. 22,000 N

OpenStudy (anonymous):

sorry i dont know i m a cs student

OpenStudy (anonymous):

there is a physics group for it

OpenStudy (anonymous):

the answer is C. 22,000 N because Pressure = Force/Area where pressure is in pascal, force is in newton and area is in meter square.

OpenStudy (anonymous):

Jaskaran's got it \[F_p = P*A\] \[F_p = 400*55 = 22000N\]